OpenStudy (anonymous):

What is the inverse function of y=4*(9)^x If I solve for x what's my answer?

OpenStudy (anonymous):

No I have no idea on how to do this

OpenStudy (freckles):

Why don't you just compare your equation to my equation and just replace my b with your b and my a with your a then?

OpenStudy (freckles):

And then you don't have to do any work.

OpenStudy (ybarrap):

$$ y=4\times 9^x\\ \ln \cfrac{y}{4}=x\ln 9\\ $$ See what I did? Solve for x then exchange y for x  Make sense?